How do I determine whether the point (-2, -2) lies on the line 2x + 3y = 10?
Answer by mathchemprofessor(65)

If a given point satisfies the given equation of a line, then the point lies on the line. Otherwise, it is not.
Now, substitute the point (-2,-2) in 2x+3y=10, we get,
2(-2)+3(-2) =-4-6=-10≠10.
Note:≠means not equal to.
So, the point (-2,-2) does not lie on the line 2x+3y=10